OpenStudy (brinazarski):

Find the domain of f. f(x) = 4x/(16x2 + 16x − 5)

OpenStudy (brinazarski):

Can someone please tell me how to do this?

OpenStudy (anonymous):

what do u mean by domain of the function?

OpenStudy (anonymous):

first find values of x for which f(x) is undefined

OpenStudy (anonymous):

HINT: f(x) is undefined when denominator equals to zero

OpenStudy (brinazarski):

1/16?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

you need to solve this equation to find the values of x for which f(x) is undefined. (16x2 + 16x − 5)=0

OpenStudy (anonymous):

can you solve it?

OpenStudy (brinazarski):

Well, I tried to and got 1/16th, but I could try again... what i did was 16x2 + 16x = 5

OpenStudy (brinazarski):

x2+x = 5/16

OpenStudy (brinazarski):

x2+x+(1/4)=9/16

OpenStudy (brinazarski):

x+1/2 = 9/16

OpenStudy (brinazarski):

x=1/16

OpenStudy (brinazarski):

Where did I go wrong, because i plugged in 1/16 and I see that it doesn't work...

OpenStudy (anonymous):

x^2+x+(1/4)=9/6 u did wrong after this it should be (x+1/2)^2 = 9/16

OpenStudy (brinazarski):

... sqrt of 9/16?

OpenStudy (brinazarski):

1/4

OpenStudy (brinazarski):

Wow i'm stupid, thank you. What do I do next?

OpenStudy (anonymous):

3/4 and -3/4

OpenStudy (brinazarski):

No, i mean 1/4 is when it's undefined

OpenStudy (anonymous):

so domain of f(x) is x belongs to all real numbers where x is not equal to 3/4 and -3/4

OpenStudy (brinazarski):

And 1/4? Where'd you get 3/4?

OpenStudy (anonymous):

square root of 9/16 is 3/4 and -3/4

OpenStudy (brinazarski):

Yes, but then you have x+1/2 = 3/4 (and I guess the negative as well)

OpenStudy (brinazarski):

So x = 1/4

OpenStudy (anonymous):

oh yes

OpenStudy (brinazarski):

or -5/4?

OpenStudy (anonymous):

so, x=1/4 and -5/4

OpenStudy (brinazarski):

The -5/4 doesn't seem to work, so disregard it?

OpenStudy (anonymous):

so domain of f(x) is x belongs to all real numbers where x is not equal to 1/4 and -5/4

OpenStudy (brinazarski):

I plugged it in and it doesn't seem to equal to 0

OpenStudy (brinazarski):

wait

OpenStudy (brinazarski):

nvm

OpenStudy (brinazarski):

yes it does

OpenStudy (anonymous):

got it?

OpenStudy (brinazarski):

Mhmm! Thank you so much!